find the sum of the first 20 positive multiples of 3? this is algebera 3
Answer by bam878s(77)

the first 20 multiples of 3 are:
3, 6, 9, 12, 15, etc
we need the sum
3 + 6 + 9 + 12 + 15 + ....
Note 3 = 3*1
6 = 3*2
9 = 3*3
12 = 3*4
15 = 3*5
etc.
So, we need:
3*1 + 3*2 + 3*3 + 3*4 + 3*5 + ... +3*20
= 3[1 + 2 + 3 + 4 + 5 + ... + 20]
= 3[210]
= 630